OpenStudy (anonymous):

Took a test and missed this. How do you solve? x/4-3 = x/3 +4

OpenStudy (anonymous):

Get both x values on the same side of the equation: x/4 - x/3 -3 = 4 Combine them, with a common denominator: 3x/12 - 4x/12 -3 =4 -x/12 -3 =4 Combine the constants: -x/12 = 7 Now multiply both sides by -12 x = 7(-12) = -84

OpenStudy (anonymous):

Thanks, I see it now. so simple.